the following table. Cutting the plant is operated at full capacity? X dozen X dozen
Answer #1

let sleeveless = x

short sleeve = y

long sleeve = z

setting up the equations

9x + 12y + 15z = 93

22x + 24y + 28z = 188

6x + 8y + 8z = 56

Step 1: Multiply first equation by and add the result to the second equation. The result is 9x+ 12y 15 z 26 - 93 118 16 3 - 56 Step 2: Multiply first equation by - 2 and add the result to the third equation. The result is 93 118 9x+ 12y 15 z 26 16 3 6 Step 3: Get rid of the fractions by multiplying second equation by 3. After multiplying we have the following system - 93 16 y 26 z 118 Step 4: solve for z 2-6 Step 5: solve for y 16y - 26z -118 16y - 26 3--118

sleeveless = 2 dozens

short sleeve = 2.5 dozens

long sleeve = 3 dozens